Shift report 20000831 08:00-16:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Thu Aug 31 2000 - 15:57:59 EDT

  • Next message: http account: "Shift report 20000831 16:00-24:00"

    ------------------------------------------------------------
    Date: Thu Aug 31 15:57:58 2000, Shift: 08:00-16:00
    Supervisor: ZM
    , with: EJK
    Log book#5 pages 64 - 65
    
    Shift Summary:
    ------------------------------------------------------------
    1. Machine stoped at 9:23 and no beam up to end of
       the shift.
    
    2. Run#2484 collected 183 K collision events.
    
    3. FV and ZM got controled access and:
       (i)  Recycled MRS HV maiframe.
       (ii) Tiles have been removed.
       MRS moved to 40 deg.; FFS and BFS moved to 4 deg.
       Problem to set D2 on. 
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Thu Aug 31 2000 - 15:58:15 EDT